# Spectral finiteness, quantum norm continuity and classical points

### 13 Mar 2026 v2

@@ Introduction
Classically, the two properties (norm continuity and spectral finiteness) are indeed equivalent: [15, p.257] addresses this for connected second-countable groups, [27, Corollary 2] in the broader setup of arbitrary compact groups represented norm-continuously on Banach spaces, and [8, Theorem 3.10] provides a number of equivalent characterizations.
@@
There are alternative possible notions of norm continuity for compact-quantum-group representations, beyond those featuring in Theorem˜0.1. [8, Definition 3.6(3)], for instance, expands the scope of the discussion to representations on Banach (as opposed to Hilbert) spaces $E$ in the sense of [8, Definition 3.1]: coassociative maps $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}E$ into the injective [12, §1.1] Banach-space tensor product satisfying the customary density condition (see ˜1-6 below). In that Banach-space context, said norm (or uniform) continuity is defined via [8, Lemma 3.7] as
@@
For a compact quantum group $\mathbb{G}$} represented via $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}E$ on a Banach space $E$ the following are equivalent.
@@
This reverses the implication recorded as [8, Lemma 3.8], but note that the hypothesis is arguably somewhat excessive: classically, one need not require the full power of ˜0-1, for in that case norm uniformity translates directly to the map in question being continuous when restricted to the space of pure states of ${\mathcal{C}}({\mathbb{G}})$ (i.e. $\mathbb{G}$} itself, with points identified with their respective Dirac-delta measures). This latter remark too has a quantum analogue, in the form of Theorem˜0.3 below. To make sense of the statement:
@@
    If the Pontryagin dual [29, Theorem 3.3.2] $\Gamma:=\widehat{{\mathbb{G}}}$ has tempered decay in the sense that
    $\begin{gathered}\exists\left(C>0\right)\forall\left(\alpha\in\operatorname{Irr}({\mathbb{G}})\right)\left(\left\|u^{\alpha}\right\|_{\infty}^{{\mathcal{C}}({\mathbb{G}})}>C\right)\\ \left\|u^{\alpha}\right\|_{\infty}^{{\mathcal{C}}({\mathbb{G}})}:=\sup_{\begin{subarray}{c}v,w\in{\mathbb{C}}^{\dim\alpha}\\ \|v\|,\|w\|\leq 1\end{subarray}}\left\|\Braket{v\mid u^{\alpha}\mid w}:=\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{i}}w_{j}\right\|.\end{gathered}$
    $U$ is uniform≤1 if and only if it has finite spectrum.
@@
    In particular, said equivalence holds if $\mathbb{G}$} has a classical point, in the sense that ${\mathcal{C}}({\mathbb{G}})$ has a multiplicative state.
@@ Acknowledgments
Initial drafts have benefited from valuable input from M. Brannan.
@@ 1 Quantum uniform continuity in several guises, and consequences
Some of the usual operator-algebraic notation and conventions will be operative: $A\bullet B$ with $\bullet\in\{\otimes,\ \underline{\otimes},\ \overline$$\otimes$}\} stand for the ordinary algebraic, and spatial (or minimal or injective) [28, Definitions IV.4.8 and IV.5.1] $C^{*}$ and $W^{*}$ tensor products respectively (between operator algebras of context-appropriate type), M()M(-) is the multiplier algebra [2, Theorem II.7.3.1], ${\mathcal{L}}(-,-)$ and ${\mathcal{K}}(-,-)$ denote spaces of bounded and compact operators respectively between Hilbert/Banach spaces, and we refer freely to slice maps ([2, §§II.9.7.1, III.2.2.6], [19, Lemma 1.5], [30, §1], etc.)
@@
For compact quantum $\mathbb{G}$} with underlying function algebra ${\mathcal{C}}({\mathbb{G}})$ we work throughout with Woronowicz’s matrix coefficients ([34, §§1 and 4], [25, §1, pp.2-3], [23, §§1.3 and 1.4], etc.)
@@
spanning the (unique dense [18, Theorem 3.1.7]) Hopf *-algebra of representative functions on a compact quantum group $\mathbb{G}$}. Unitary $\mathbb{G}$}-representations on Hilbert spaces $H$ are as in [11, §2.2]:
@@
diagonal. One effect this will have is to render the $u^{\alpha}_{ij}$, (via [34, (5.14) and (5.15)] or [23, Theorem 1.4.3]), mutually orthogonal with respect to both versions of the inner product induced by the Haar state $h:=h_{{\mathbb{G}}}$ of [34, §4] or [23, §1.2]:
@@
the index indicating the position (left/right) of the product’s linear argument.
@@
The extent to which this last equality holds formally, with the right-hand side norm-convergent, is the focus of amply studied Fourier decay-at-infinity phenomena: quantum analogues of the familiar [4, §I.2] Riemann-Lebesgue lemma in classical Fourier analysis, as alluded to briefly in [1, §1, p.34], say. We revisit the matter in the course of the proof of Theorem˜0.1.
@@
The equivalence ˜a $\Leftrightarrow$ ˜b is covered by [8, Theorem 3.16], while ˜a is plainly stronger than the other conditions: finite spectrum in fact entails
@@
It remains to verify ˜c $\Rightarrow$ ˜a, which task the proof henceforth turns to. Reprising the notation of ˜1-4 for the usual conditional expectation [2, Definition II.6.10.1] ${\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{L}}(H)\xrightarrow[]{E}\mathrel{\mkern-14.0mu}\rightarrow{\mathcal{L}}(H)$, Proposition˜1.2 below ensures that that equation in fact holds as written, with a norm-convergent right-hand side. Applied to $x:=U$, ˜1-4 reads
@@
are the respective projections (introduced in [25, Theorem 1.5] as $E^{\alpha}$) onto the $\alpha$-isotypic components (cf. [3, remarks preceding Proposition 13]) of the representation $U$. Said norm convergence of mutually orthogonal projections, naturally, requires that only finitely many be non-zero. $\blacksquare$
The proof of the auxiliary Proposition˜1.2 below will precisely parallel its analogue [26, Lemma 5.2] in the context of reduced crossed products [6, Definition 4.1.4] resulting from actions by countable (classical, discrete) groups on $C^{*}$-algebras (so in our language, that result’s $\mathbb{G}$} would be an abelian, i.e. dual-classical, compact quantum group). That source leverages weaker convergence into its stronger, norm counterpart via a device credited there to G. Elliott and U. Haagerup and reliant on Dini’s [14, Theorem 12.1] [^1] Said invocation of Dini, typically stated for sequences of real-valued functions on compact Hausdorff spaces, appears to be the one aspect of the proof of [26, Lemma 5.2] relying in any substantive fashion on any countability hypotheses. We will see in the course of the proof of Proposition 1.2 that such countability assumptions can be dispensed with. (see also [20, Lemma 3.1] for essentially the same gadget).
@@
in the $\sigma$-strong topology of [28, Definition II.2.3] (the strongest of the locally convex sub-norm topologies examined in [28, §II.2]): the individual summands $\frac{q_{\alpha}}{F_{\alpha jj}}T^{\alpha}_{ij}T^{\alpha*}_{ij}$ are (by [34, (5.14)] again) mutually orthogonal projections whose ranges exhaust the underlying Hilbert space.
@@
One again has spectral idempotents $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})\circ\rho\in{\mathcal{L}}(E)$ analogous to ˜1-5 and, by [8, Proposition], the more familiar unitary representations $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ can be recast as representations on the underlying Banach spaces by
(1-7) $H\ni v\xmapsto{\quad}U^{*}(1\otimes v)\in{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}H.$
@@
While [8, Lemma 3.8] notes only that having finite spectrum implies ˜0-1, the converse is also straightforward: weak-to-norm continuity is an extremely strong condition to impose, holding only in trivial circumstances.
@@
For classical $\mathbb{G}$} bounded and plain uniformity are equivalent: as noted in the Introduction following Theorem˜0.2, uniformity≤1 already implies spectrum finiteness and hence [8, Lemma 3.8] also ordinary uniformity. It is not inapposite, then, to ask whether the equivalence survives quantization.
As further motivation and surrounding context for Theorem˜0.3, recall [7, Remark 1.3] to the effect that the classical [27, Corollary 2] characterization of norm-continuous compact-group representations as precisely those with finite spectra is recovered by non-classical means from [7, Theorem 0.1] for compact quantum groups $\mathbb{G}$} with a classical point (i.e. with ${\mathcal{C}}({\mathbb{G}})$ admitting a multiplicative state); by extension, the same goes for Theorem˜0.1 (which removes the classical-point constraint). Theorem˜0.3 describes how, upon assuming a small amount of classical behavior, said classical result replicates in terms of what is perhaps the most straightforward analogue of quantum uniform continuity for a unitary $\mathbb{G}$}-representation.
@@
The device isolated by Proposition˜1.6 below, implicitly also driving the proof of [30, Proposition 2.2], will allow some symmetrization in the condition on $U$ imposed by Lemma˜1.4. Some vocabulary will help compress the statement.
@@
    As in ˜a, for nets of functionals of the form
    $M\ni T\xmapsto{\quad\omega_{v,w}\quad}\Braket{v\mid Tw}\in{\mathbb{C}}$
    for $v,w\in H$ in a Hilbert space on which $M$ is represented faithfully as $M\leq{\mathcal{L}}(H)$.
@@
Note that ˜a concerns all of $M^{*}$, whereas the other $\bullet_{\ell}$ specialize to (at least) weak-continuous functionals. The proof of [30, Proposition 2.2], in present language, shows effectively that the strongest left-handed statement ˜a is implied by its right-handed counterpart phrased for nets $(\psi_{\lambda})_{\lambda}\subset N_{*,\leq 1}$ (unit ball of the predual of $N$) assumed convergent $N^{*}_{\leq 1}$. This is easily recast as ˜br though: the net in question is certainly Cauchy and hence $(\varphi_{\lambda}-\varphi_{\lambda^{\prime}})_{\lambda,\lambda^{\prime}}$ is 0-convergent; consequently,
@@
˜c $\Rightarrow$ ˜b by weak-in-weak norm density [28, Theorem II.2.6(iii)].
In order to complete the proof of the proposition as a whole by confirming that ˜er implies ˜a, it remains to observe that the aforementioned net $(\psi_{\lambda})_{\lambda}\subset N_{*,\leq 1}$ can be chosen so as to range over any set ${\mathcal{S}}\subset N_{\leq 1}^{*}$ detecting norms in $N$ in the following sense:
$\forall\left(T\in N\right)\forall\left(\varepsilon>0\right)\exists\left(\varphi\in{\mathcal{S}}\right)\left(|\varphi T|\geq\|T\|-\varepsilon\right).$
@@
The aforementioned symmetrization afforded by Proposition˜1.6 refers to switching Lemma˜1.4’s criterion to $U$’s right-hand leg. For $v\in H$ we write
Corollary 1.7.
A unitary representation $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ of a compact quantum group $\mathbb{G}$} is uniform≤1 as a Banach-space representation if and only if
Proof.
Lemma˜1.4 and Proposition˜1.6’s equivalence ˜a $\Leftrightarrow$ ˜dr equate uniformity≤1 with the weak-norm convergence of
$\omega_{v,w}\xmapsto{\quad}\left(\operatorname{id}\otimes\omega_{v,w}\right)U\in{\mathcal{L}}(H),\quad v,w\in H_{\leq 1}.$
We may thus assume weakly convergent nets
$\left(v_{\lambda}\right)_{\lambda}\xrightarrow[\quad\lambda\quad]{\quad}v\quad\text{and}\quad\left(w_{\lambda}\right)_{\lambda}\xrightarrow[\quad\lambda\quad]{\quad}w$
in the unit ball $H_{\leq 1}$, whence $\omega_{v_{\lambda},w_{\lambda}}\to\omega_{v,w}$. Approximate $v$ and $w$ arbitrarily well by finitely supported vectors $\overline{v}$ and $\overline{w}$ respectively, assuming also that
$\operatorname{\mathrm{supp}}\left(v_{\lambda}-\overline{v}\right)\to\infty\quad\text{and}\quad\operatorname{\mathrm{supp}}\left(w_{\lambda}-\overline{w}\right)\to\infty.$
Because finite truncations
$\sum_{\alpha\in{\mathcal{F}}}\left(1\otimes P^{\alpha}\right)U,\quad\text{finite }{\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$
belong to the algebraic tensor product ${\mathcal{O}}({\mathbb{G}})\otimes{\mathcal{L}}(H)$, the norm convergence of the truncated
$\left(\operatorname{id}\otimes\omega_{\bullet}\right)U,\quad\bullet\in\left\{(\overline{v},\overline{w}),\ \left(v_{\lambda}-\overline{v},\overline{w}\right),\ \left(\overline{v},w_{\lambda}-\overline{w}\right)\right\}$
is not at issue. This confirms the equivalence
$\left(\operatorname{id}\otimes\omega_{v_{\lambda},w_{\lambda}}\right)U\xrightarrow[\quad\lambda\quad]{\quad\text{norm}\quad}\left(\operatorname{id}\otimes\omega_{v,w}\right)U\quad\xLeftrightarrow{\quad}\quad\left(\operatorname{id}\otimes\omega_{v_{\lambda}-\overline{v},w_{\lambda}-\overline{w}}\right)U\xrightarrow[\quad\lambda\quad]{\quad\text{norm}\quad}0,$
finishing the proof. $\blacksquare$
@@
  1. ˜1 Assume infinitely many spectral projections $P^{\alpha}$ of $U$ non-zero, entailing the corresponding non-vanishing Fourier coefficients $x^{\alpha}_{ij}$, $1\leq i,j\leq\dim\alpha$ of $x:=U$ defined by ˜1-3 (having fixed the $u^{\alpha}$-engendering bases of the $\alpha$ as explained preceding ˜1-2, so as to ensure the $u^{\alpha}_{ij}$ are mutually orthogonal). The hypothesis ensures the existence of norm-1
    $v_{0}=\left(v_{0i}\right)_{1\leq i\leq\dim\alpha}\quad\text{and}\quad w_{0}=\left(w_{0i}\right)_{1\leq i\leq\dim\alpha}\quad:\quad\left\|\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{0i}}w_{0j}\right\|>C.$
    To procure suitable $\omega_{v,w}$ amenable to Corollary˜1.7, set
    $v:=\sum_{i=1}^{\dim\alpha}v_{0i}e_{i}\quad\text{and}\quad w:=\sum_{i=1}^{\dim\alpha}w_{0i}e_{i},\quad e_{i}\in\operatorname{\mathrm{im}}x^{\alpha}_{ii}\leq\operatorname{\mathrm{im}}P^{\alpha}\text{ orthonormal}.$
    We have
    $\left\|(\operatorname{id}\otimes\omega_{v,w})U\right\|=\left\|\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{0i}}w_{0j}\right\|>C,$
    affording the desired conclusion that uniformity≤1 must fail.
  2. ˜2 $\mathbb{G}$} has a classical point precisely when the counit $\varepsilon$ has a continuous extension to all of ${\mathcal{C}}({\mathbb{G}})$ (see [10, paragraph preceding Theorem 2.2]; also Lemma˜1.8 below for additional equivalent characterizations), whence a state assigning the value 11 to every $u^{\alpha}_{ii}$. Those elements, then, must all have norm 1.
@@
In reference to the property featuring in [7, Theorem 0.1] of compact quantum groups having at least one classical point (one version of the above-mentioned “partial classical behavior”), Lemma˜1.8 records a simple remark for whatever future-reference purpose it may serve. We will be referring to the usual convolution product [21, Definition 1.4.1], denoted here simply by $\cdot$” or juxtaposition, and defined as the precomposition
@@
on the algebraic dual of a coalgebra. Note also that it (or rather its appropriate extension) makes the state space ${\mathcal{S}}({\mathcal{C}}({\mathbb{G}}))$ of the $C^{*}$-algebra ${\mathcal{C}}({\mathbb{G}})$ into a compact Hausdorff [9, Theorem V.3.1] semigroup under its weak topology. The same goes for the unit ball ${\mathcal{C}}({\mathbb{G}})^{*}_{\leq 1}$ in the continuous dual of ${\mathcal{C}}({\mathbb{G}})$: the norm dual to $\underline{\otimes}$ being cross [32, Proposition T.3.11],

### 20 Mar 2026 v3

@@ Introduction
Classically, the two properties (norm continuity and spectral finiteness) are indeed equivalent: [15, p.257] addresses this for connected second-countable groups, [28, Corollary 2] in the broader setup of arbitrary compact groups represented norm-continuously on Banach spaces, and [9, Theorem 3.10] provides a number of equivalent characterizations.
@@
There are alternative possible notions of norm continuity for compact-quantum-group representations, beyond those featuring in Theorem˜0.1. [9, Definition 3.6(3)], for instance, expands the scope of the discussion to representations on Banach (as opposed to Hilbert) spaces $E$ in the sense of [9, Definition 3.1]: coassociative maps $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\varepsilon}E$ into the injective [27, §3.1] Banach-space tensor product satisfying the customary density condition (see ˜1-6 below). In that Banach-space context, said norm (or uniform) continuity is defined via [9, Lemma 3.7] as
@@
For a compact quantum group $\mathbb{G}$} represented via $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\varepsilon}E$ on a Banach space $E$ the following are equivalent.
@@
This reverses the implication recorded as [9, Lemma 3.8], but note that the hypothesis is arguably somewhat excessive: classically, one need not require the full power of ˜0-1, for in that case norm uniformity translates directly to the map in question being continuous when restricted to the space of pure states of ${\mathcal{C}}({\mathbb{G}})$ (i.e. $\mathbb{G}$} itself, with points identified with their respective Dirac-delta measures). This latter remark too has a quantum analogue, in the form of Theorem˜0.3 below. To make sense of the statement:
@@
    If the Pontryagin dual [30, Theorem 3.3.2] $\Gamma:=\widehat{{\mathbb{G}}}$ has tempered decay in the sense that
    $\begin{gathered}\exists\left(C>0\right)\forall\left(\alpha\in\operatorname{Irr}({\mathbb{G}})\right)\left(\left\|u^{\alpha}\right\|_{\infty}^{{\mathcal{C}}_{r}({\mathbb{G}})}>C\right)\\ \left\|u^{\alpha}\right\|_{\infty}^{{\mathcal{C}}({\mathbb{G}})}:=\sup_{\begin{subarray}{c}v,w\in{\mathbb{C}}^{\dim\alpha}\\ \|v\|,\|w\|\leq 1\end{subarray}}\left\|\Braket{v\mid u^{\alpha}\mid w}:=\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{i}}w_{j}\right\|,\end{gathered}$
    then $U$ is uniform≤1 if and only if it has finite spectrum.
@@
    In particular, said equivalence holds if the reduced version of $\mathbb{G}$} has a classical point, in the sense that ${\mathcal{C}}_{r}({\mathbb{G}})$ has a multiplicative state. Equivalently: if $\mathbb{G}$} is coamenable [1, Definition 3.1].
@@ Acknowledgments
Initial drafts have benefited from M. Brannan’s valuable input
@@ 1 Quantum uniform continuity in several guises, and consequences
Some of the usual operator-algebraic notation and conventions will be operative: $A\bullet B$ with $\bullet\in\{\otimes,\ \underline{\otimes},\ \overline$$\otimes$}\} stand for the ordinary algebraic, and spatial (or minimal or injective) [29, Definitions IV.4.8 and IV.5.1] $C^{*}$ and $W^{*}$ tensor products respectively (between operator algebras of context-appropriate type), M()M(-) is the multiplier algebra [3, Theorem II.7.3.1], ${\mathcal{L}}(-,-)$ and ${\mathcal{K}}(-,-)$ denote spaces of bounded and compact operators respectively between Hilbert/Banach spaces, and we refer freely to slice maps ([3, §§II.9.7.1, III.2.2.6], [19, Lemma 1.5], [31, §1], etc.)
@@
For compact quantum $\mathbb{G}$} with underlying function algebra ${\mathcal{C}}({\mathbb{G}})$ we work throughout with Woronowicz’s matrix coefficients ([35, §§1 and 4], [25, §1, pp.2-3], [23, §§1.3 and 1.4], etc.)
@@
spanning the (unique dense [18, Theorem 3.1.7]) Hopf *-algebra of representative functions on a compact quantum group $\mathbb{G}$}. Unitary $\mathbb{G}$}-representations on Hilbert spaces $H$ are as in [12, §2.2]:
@@
diagonal. One effect this will have is to render the $u^{\alpha}_{ij}$, (via [35, (5.14) and (5.15)] or [23, Theorem 1.4.3]), mutually orthogonal with respect to both versions of the inner product induced by the Haar state $h:=h_{{\mathbb{G}}}$ of [35, §4] or [23, §1.2]:
@@
the index indicating the position (left/right) of the product’s linear argument. ${\mathcal{C}}_{r}({\mathbb{G}})$ denotes the reduced version of $\mathcal{G}$}, i.e. [23, post Corollary 1.7.5] the image of ${\mathcal{C}}({\mathbb{G}})$ through the GNS representation attached to $h$.
@@
The extent to which this last equality holds formally, with the right-hand side norm-convergent, is the focus of amply studied Fourier decay-at-infinity phenomena: quantum analogues of the familiar [5, §I.2] Riemann-Lebesgue lemma in classical Fourier analysis, as alluded to briefly in [2, §1, p.34], say. We revisit the matter in the course of the proof of Theorem˜0.1.
@@
The equivalence ˜a $\Leftrightarrow$ ˜b is covered by [9, Theorem 3.16], while ˜a is plainly stronger than the other conditions: finite spectrum in fact entails
@@
It remains to verify ˜c $\Rightarrow$ ˜a, which task the proof henceforth turns to. Reprising the notation of ˜1-4 for the usual conditional expectation [3, Definition II.6.10.1] ${\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{L}}(H)\xrightarrow[]{E}\mathrel{\mkern-14.0mu}\rightarrow{\mathcal{L}}(H)$, Proposition˜1.2 below ensures that that equation in fact holds as written, with a norm-convergent right-hand side. Applied to $x:=U$, ˜1-4 reads
@@
are the respective projections (introduced in [25, Theorem 1.5] as $E^{\alpha}$) onto the $\alpha$-isotypic components (cf. [4, remarks preceding Proposition 13]) of the representation $U$. Said norm convergence of mutually orthogonal projections, naturally, requires that only finitely many be non-zero. $\blacksquare$
The proof of the auxiliary Proposition˜1.2 below will precisely parallel its analogue [26, Lemma 5.2] in the context of reduced crossed products [7, Definition 4.1.4] resulting from actions by countable (classical, discrete) groups on $C^{*}$-algebras (so in our language, that result’s $\mathbb{G}$} would be an abelian, i.e. dual-classical, compact quantum group). That source leverages weaker convergence into its stronger, norm counterpart via a device credited there to G. Elliott and U. Haagerup and reliant on Dini’s [14, Theorem 12.1] [^1] Said invocation of Dini, typically stated for sequences of real-valued functions on compact Hausdorff spaces, appears to be the one aspect of the proof of [26, Lemma 5.2] relying in any substantive fashion on any countability hypotheses. We will see in the course of the proof of Proposition 1.2 that such countability assumptions can be dispensed with. (see also [20, Lemma 3.1] for essentially the same gadget).
@@
in the $\sigma$-strong topology of [29, Definition II.2.3] (the strongest of the locally convex sub-norm topologies examined in [29, §II.2]): the individual summands $\frac{q_{\alpha}}{F_{\alpha jj}}T^{\alpha}_{ij}T^{\alpha*}_{ij}$ are (by [35, (5.14)] again) mutually orthogonal projections whose ranges exhaust the underlying Hilbert space.
@@
One again has spectral idempotents $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})\circ\rho\in{\mathcal{L}}(E)$ analogous to ˜1-5 and, by [9, Proposition], the more familiar unitary representations $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ can be recast as representations on the underlying Banach spaces by
(1-7) $H\ni v\xmapsto{\quad}U^{*}(1\otimes v)\in{\mathcal{C}}({\mathbb{G}})\otimes_{\varepsilon}H.$
@@
While [9, Lemma 3.8] notes only that having finite spectrum implies ˜0-1, the converse is also straightforward: weak-to-norm continuity is an extremely strong condition to impose, holding only in trivial circumstances.
@@
For classical $\mathbb{G}$} bounded and plain uniformity are equivalent: as noted in the Introduction following Theorem˜0.2, uniformity≤1 already implies spectrum finiteness and hence [9, Lemma 3.8] also ordinary uniformity. It is not inapposite, then, to ask whether the equivalence survives quantization.
As further motivation and surrounding context for Theorem˜0.3, recall [8, Remark 1.3] to the effect that the classical [28, Corollary 2] characterization of norm-continuous compact-group representations as precisely those with finite spectra is recovered by non-classical means from [8, Theorem 0.1] for compact quantum groups $\mathbb{G}$} with a classical point (i.e. with ${\mathcal{C}}({\mathbb{G}})$ admitting a multiplicative state); by extension, the same goes for Theorem˜0.1 (which removes the classical-point constraint). Theorem˜0.3 describes how, upon assuming a small amount of classical behavior, said classical result replicates in terms of what is perhaps the most straightforward analogue of quantum uniform continuity for a unitary $\mathbb{G}$}-representation.
@@
The device isolated by Proposition˜1.6 below, implicitly also driving the proof of [31, Proposition 2.2], will allow some symmetrization in the condition on $U$ imposed by Lemma˜1.4. Some vocabulary will help compress the statement.
@@
    Setting $\omega_{v,w}:=\Braket{v\mid\bullet w}\in M_{*}$ for $v,w\in H$ in a Hilbert space on which $M$ is represented faithfully as $M\leq{\mathcal{L}}(H)$,
    $\omega_{v,w}\xmapsto{\quad}(\omega_{v,w}\otimes\operatorname{id})x\in{\mathcal{C}}({\mathbb{G}})$
    turns weak-Cauchy nets into their norm counterparts.
@@
Note that ˜a concerns all of $M^{*}$, whereas the other $\bullet_{\ell}$ specialize to (at least) weak-continuous functionals. The proof of [31, Proposition 2.2], in present language, shows effectively that the strongest left-handed statement ˜a is implied by its right-handed counterpart phrased for nets $(\psi_{\lambda})_{\lambda}\subset N_{*,\leq 1}$ (unit ball of the predual of $N$) assumed convergent $N^{*}_{\leq 1}$. This is easily recast as ˜br though: the net in question is certainly Cauchy and hence $(\psi_{\lambda}-\psi_{\lambda^{\prime}})_{\lambda,\lambda^{\prime}}$ is 0-convergent; consequently,
@@
˜c $\Rightarrow$ ˜b by weak-in-weak norm density [29, Theorem II.2.6(iii)].
In order to complete the proof of the proposition as a whole by confirming that ˜dr implies ˜a, it remains to observe that the aforementioned net $(\psi_{\lambda})_{\lambda}\subset N_{*,\leq 1}$ can be chosen so as to range over any set ${\mathcal{S}}\subset N_{\leq 1}^{*}$ detecting norms in $N$ in the following sense:
$\forall\left(T\in N\right)\forall\left(\varepsilon>0\right)\exists\left(\psi\in{\mathcal{S}}\right)\left(|\psi T|\geq\|T\|-\varepsilon\right).$
@@
The aforementioned symmetrization afforded by Proposition˜1.6 refers to switching Lemma˜1.4’s criterion to $U$’s right-hand leg, as will become apparent presently.
@@
  1. ˜1 Assume infinitely many spectral projections $P^{\alpha}$ of $U$ non-zero, entailing the corresponding non-vanishing Fourier coefficients $x^{\alpha}_{ij}$, $1\leq i,j\leq\dim\alpha$ of $x:=U$ defined by ˜1-3 (having fixed the $u^{\alpha}$-engendering bases of the $\alpha$ as explained preceding ˜1-2, so as to ensure the $u^{\alpha}_{ij}$ are mutually orthogonal). The tempered-decay hypothesis ensures the existence of norm-1
    $v_{\alpha 0}=\left(v_{\alpha 0i}\right)_{1\leq i\leq\dim\alpha}\quad\text{and}\quad w_{\alpha 0}=\left(w_{\alpha 0i}\right)_{1\leq i\leq\dim\alpha}\quad:\quad\left\|\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{\alpha 0i}}w_{\alpha 0j}\right\|>C.$
    Now consider $\omega_{\alpha}:=\omega_{v_{\alpha},w_{\alpha}}:=\Braket{v_{\alpha}\mid\bullet w_{\alpha}}\in{\mathcal{L}}(H)_{*}$ for
    $v_{\alpha}:=\sum_{i=1}^{\dim\alpha}v_{\alpha 0i}e_{i}\quad\text{and}\quad w_{\alpha}:=\sum_{i=1}^{\dim\alpha}w_{\alpha 0i}e_{i},\quad e_{i}\in\operatorname{\mathrm{im}}x^{\alpha}_{ii}\leq\operatorname{\mathrm{im}}P^{\alpha}\text{ orthonormal},$
    so that by construction we have
    $\left\|(\operatorname{id}\otimes\omega_{\alpha})U\right\|=\left\|\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{\alpha 0i}}w_{\alpha 0j}\right\|>C.$
    By Alaoglu’s [10, Theorem V.3.1], the bounded family $\{\omega_{\alpha}\}_{\alpha}$ supports a weak-Cauchy net which we index by the $\alpha$ themselves, with the understanding that $\alpha$ eventually avoids every finite subset of $\operatorname{Irr}({\mathbb{G}})$.
    Now, on the one hand the net $\left((\operatorname{id}\otimes\omega_{\alpha})U\right)_{\alpha}$ must norm-converge in ${\mathcal{C}}_{r}({\mathbb{G}})$ to an element of norm $\geq C>0$. On the other, matrix-coefficient Haar orthogonality forces
[…]
    $h$ being faithful [23, Corollary 1.7.5] on ${\mathcal{C}}_{r}({\mathbb{G}})$, the desired conclusion that uniformity≤1 must fail follows from that property’s characterization as Proposition˜1.6’s right-handed ˜d.
  2. ˜2 $\mathbb{G}$} has a classical point precisely when the counit $\varepsilon$ has a continuous extension to all of ${\mathcal{C}}({\mathbb{G}})$ (see [11, paragraph preceding Theorem 2.2]; also Lemma˜1.7 below for additional equivalent characterizations), whence a state on ${\mathcal{C}}_{r}({\mathbb{G}})$ assigning the value $1$ to every $u^{\alpha}_{ii}$. Those elements, then, must all have norm 1 and the conclusion follows from part ˜1.
@@
In reference to the property featuring in [8, Theorem 0.1] of compact quantum groups having at least one classical point (one version of the above-mentioned “partial classical behavior”), Lemma˜1.7 records a simple remark for whatever future-reference purpose it may serve. We will be referring to the usual convolution product [21, Definition 1.4.1], denoted here simply by $\cdot$” or juxtaposition, and defined as the precomposition
@@
on the algebraic dual of a coalgebra. Note also that it (or rather its appropriate extension) makes the state space ${\mathcal{S}}({\mathcal{C}}({\mathbb{G}}))$ of the $C^{*}$-algebra ${\mathcal{C}}({\mathbb{G}})$ into a compact Hausdorff [10, Theorem V.3.1] semigroup under its weak topology. The same goes for the unit ball ${\mathcal{C}}({\mathbb{G}})^{*}_{\leq 1}$ in the continuous dual of ${\mathcal{C}}({\mathbb{G}})$: the norm dual to $\underline{\otimes}$ being cross [33, Proposition T.3.11],